OpenStudy (anonymous):

The equation of a a circle is (x+3)^2 + (y-5)^2 = 81. Determine the coordinates of the center of the circle && the length of the circle A. center = (-3,5) radius = 9 B. center = (3,5) radius = 9 C. center = (-5,3) radius = 81 D. center = (5,3) radius = 81

OpenStudy (anonymous):

In the general equation of circle, (x - a)² + (y - b)² = r² Centre = (a, b) Radius = r (x + 3)² + (y - 5)² = 81 (x + 3)² + (y - 5)² = (9)² Comparing with general equation, Centre = (-3, 5) Radius = 9